OpenStudy (anonymous):

can someone tell me what a function is and how can you tell if its linear?? Thanks!

OpenStudy (ash2326):

@Chu~Luv~You2019 A function is a relation between two sets X and Y such that for each x there is a unique value of y. For example X= (2, 5) correspondingly Y= (4, 10) for x =2, y=4 and x=5, y=10 so this is a function Do you understand this?

OpenStudy (anonymous):

Yes ! Thank you!

OpenStudy (amistre64):

linear and affine functions are related, but a strict definition of linear would rule out affine function

OpenStudy (amistre64):

y = kx + c is affine if c is not equal to zero it is linear of c IS eqiual to zero

OpenStudy (amistre64):

a function is linear if: f(x+y) = f(x) + f(y) and f(kx) = k f(x)

OpenStudy (ash2326):

ha ha, so OK, I'll explain linear suppose we have function \[y=2x\] notice the variation in x and y x= 1, 2, 3 y=2, 4, 6 As x increases by 1 , y increases by 2 There is a definite variation in y. Do you get this part? @Chu~Luv~You2019

OpenStudy (anonymous):

yes i believe i do .. ^^ ~ ~ ~